1.     A transportation problem involves the following costs, supply, and demand.

 

 

To

 

 

From

1

2

3

4

Supply

1

$500

750

300

450

12

2

650

800

400

600

17

3

400

700

500

550

11

Demand

10

10

10

10

 

Required:

             i.           Find the initial solution using the northwest corner method, the minimum cell cost method, and Vogel's Approximation Method. Compute total cost for each.

           ii.           Using the VAM initial solution, find the optimal solution using the modified distribution method (MODI).



1
Expert's answer
2021-09-24T12:17:41-0400

i. Find the initial solution using the northwest corner method, the minimum cell cost method, and Vogel's Approximation Method. Compute total cost for each.




Northwest Corner Method




To


From 1 2 3 4 Supply


1 $500 750 300 450 12


10 2


2 650 800 400 600 17


8 9


3 400 700 500 550 11


1 10


Demand 10 10 10 10 40




Z = 500X11 + 750X12 + 800X22 + 400X23 + 500X33 + 550X34


Z = 500 (10) + 750 (2) + 800 (8) + 400 (9) + 500 (1) + 550 (10)


Z = 5,000 + 1,500 + 6,400 + 3,600 + 500 + 5,500


Answer: Total Cost: Z = $ 22,500




Minimum Cell Cost Method




To


From 1 2 3 4 Supply


1 $500 750 300 450 12


10 2


2 650 800 400 600 17


10 7


3 400 700 500 550 11


10 1


Demand 10 10 10 10 40




Z = 300X13 + 450X14 + 800X22 + 600X24 + 400X31 + 550X34


Z = 300 (10) + 450 (2) + 800 (10) + 600 (7) + 400 (10) + 550 (1)


Z = 3,000 + 900 + 8,000 + 4,200 + 4,000 + 550


Answer: Total Cost: Z = $ 20,650




Vogel’s Approximation Method




To


From 1 2 3 4 Supply


1 $500 750 300 450 12


2 10 300


2 650 800 400 600 17


7 10 200


3 400 700 500 550 11


10 1 150


Demand 10 10 10 10 40




Z = 750X12 + 450X14 + 800X22 + 400X23 + 400X31 + 700X32


Z = 750 (2) + 450 (10) + 800 (7) + 400 (10) + 400 (10) + 700 (1)


Z = 1,500 + 4,500 + 5,600 + 4,000 + 4,000 + 700


Answer: Total Cost: Z = $ 20,300




ii. Using the VAM initial solution, find the optimal solution using the modified distribution method (MODI).




SOLUTION:




X12 : u1 + v2 = 750 Let u1 = 0


X14 : u1 + v4 = 450 Let v2 = 750


X22 : u2 + v2 = 800 Let v4 = 450


X23 : u2 + v3 = 400 Let u2 = 50


X31 : u3 + v1 = 400 Let v3 = 350


X32 : u3 + v2 = 700 Let u3 = –50


v1 = 450


vi v1 = 450 v2 = 750 v3 = 350 v4 = 450


To


ui From 1 2 3 4 Supply


u1 = 0 1 $500 750 300 450 12


2 10


u2 = 50 2 650 800 400 600 17


7 10


u3 = –50 3 400 700 500 550 11


10 1


Demand 10 10 10 10 40




X11 : k11 = 500 – 0 – 450 = 50


X13 : k13 = 300 – 0 – 350 = –50


X21 : k21 = 650 – 50 – 450 = 150


X24 : k24 = 600 – 50 – 450 = 50


X33 : k33 = 500 – (–50) – 350 = 200


X34 : k34 = 550 – (–50) – 450 = 150




To


From 1 2 3 4 Supply


1 $500 - 750 + 300 450 12


2 10


2 650 + 800 - 400 600 17


7 10


3 400 700 500 550 11


10 1


Demand 10 10 10 10 40






To


From 1 2 3 4 Supply


1 $500 750 300 450 12


2 10


2 650 800 400 600 17


7 10


3 400 700 500 550 11


10 1


Demand 10 10 10 10 40


 


X13 : u1 + v3 = 300 Let u1 = 0


X14 : u1 + v4 = 450 Let v3 = 300


X22 : u2 + v2 = 800 Let v4 = 450


X23 : u2 + v3 = 400 Let u2 = 100


X31 : u3 + v1 = 400 Let v2 = 700


X32 : u3 + v2 = 700 Let u3 = 0


v1 = 400




vi v1 = 450 v2 = 750 v3 = 350 v4 = 450


To


ui From 1 2 3 4 Supply


u1 = 0 1 $500 750 300 450 12


2 10


u2 = 100 2 650 800 400 600 17


9 8


u3 = 0 3 400 700 500 550 11


10 1


Demand 10 10 10 10 40




X11 : k11 = 500 – 0 – 400 = 100


X12 : k12 = 750 – 0 – 700 = 50


X21 : k21 = 650 – 100 – 400 = 150


X24 : k24 = 600 – 100 – 450 = 100


X33 : k33 = 500 – 0 – 300 = 200


X34 : k34 = 500 – 0 – 450 = 50




Z = 300X13 + 450X14 + 800X22 + 400X23 + 400X31 + 700X32 


Z = 300 (2) + 450 (10) + 800 (9) + 400 (8) + 400 (10) + 700 (1)


Z = 600 + 4,500 + 7,200 + 3,200 + 4,000 + 700


Answer: Z = $ 20,200